OpenStudy (anonymous):

Solve for 2x^2-5x+6=0

OpenStudy (anonymous):

Factor it out

OpenStudy (anonymous):

Not factor able?

OpenStudy (anonymous):

then use quadratic formula

OpenStudy (anonymous):

\[(-b \pm \sqrt{b^{2}-4ac}) / 2a\] a = 2 b = -5 c = 6

OpenStudy (anonymous):

Workin on it...

OpenStudy (anonymous):

should be easy

OpenStudy (anonymous):

5+- square root -23/4

OpenStudy (anonymous):

\[x = 5/4 + \sqrt{23}\iota\] and \[x = 5/4 - \sqrt{23}\iota\]

OpenStudy (anonymous):

Or 5+- SR 23i /4?

OpenStudy (anonymous):

It's all over 4, so it becomes 5/4 =) Other than that yes you're completely right

OpenStudy (anonymous):

What about the -23 in the square root?

OpenStudy (anonymous):

it becomes sqrt(23)(i) Where\[\iota = \sqrt{-1}\]

OpenStudy (anonymous):

You can find an exact value for that in the notation of 5/4 +- 4.796 i

OpenStudy (anonymous):

Alrighty then